the altitude, A, of an airplane follows a linear model given by A=mt+b where t is the time after takeoff. The plan is 1200ft off the ground at 5 seconds and 3400ft off the ground at 60 seconds. How high is the ground above sea level? What is the rate of increase in altitude over the period of 0 < t < 60
Answer by Alan3354(69443)

the altitude, A, of an airplane follows a linear model given by A=mt+b where t is the time after takeoff. The plan is 1200ft off the ground at 5 seconds and 3400ft off the ground at 60 seconds. How high is the ground above sea level? What is the rate of increase in altitude over the period of 0 < t < 60
------------------------
2 points are given: (5,1200) and (60,3400)
The slope is diffA/difft = (3400 - 1200)/(60-5) = 2200/55 = 40 ft/sec
A = 40t + GL (ground level)
1200 = 40*5 + GL
Ground level = 1000 feet
-----------------
What is the rate of increase in altitude over the period of 0 < t < 60
----
Vertical speed = 40 ft/sec = 2400 ft/min